OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. L(5, 0), M(0, 5)

OpenStudy (anonymous):

first we find the slope with the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (5,0) x1 = 5 and y1 = 0 (0,5) x2 = 0 and y2 = 5 now we sub slope(m) = (5 - 0) / (0 - 5) slope(m) = 5/-5 slope(m) = -1 now we use y = mx + b slope(m) = -1 you can use either set of points, answer will come out the same (5,0) x = 5 and y = 0 now lets sub 0 = -1(5) + b 0 = -5 + b 5 = b your equation in slope intercept form is : y = -x + 5 but we want this in standard form add x to both sides x + y = 5 ===> this is standard form

OpenStudy (anonymous):

Thanks I kinda get it you explaine every well could you help me with a cupple more ?

OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through (2, -1) and parallel to a line with slope of 3/4

OpenStudy (anonymous):

a parallel line will have the same slope, so the slope we need is 3/4. y = mx + b (2,-1) x = 2 and y = -1 slope(m) = 3/4 now lets sub and solve for b -1 = 3/4(2) + b -1 = 3/2 + b -1 - 3/2 = b -2/2 - 3/2 = b -5/2 = b equation in slope intercept form is : y = 3/4x - 5/2 --- multiply by the LCD, which is 4 4y = 3x - 10 -- subtract 3x from both sides -3x + 4y = -10 -- multiply by -1 to make x positive 3x - 4y = 10 ===> standard form

OpenStudy (anonymous):

thank you :) so much